# An adaptation of InfoMap to absorbing random walks using absorption-scaled graphs

### v3

@@ 2 Background on InfoMap
$\tilde{A}$ Adjacency matrix of a graph with an
absorbing state
@@ 2.1 The map function and the standard InfoMap algorithm
(Map function) Let $P=(p_{ij})_{i,j\in\{1,\ldots,n\}}$ be a stochastic matrix, and let $\pi=(\pi_{1},\ldots,\pi_{n})^{\operatorname{\,T}}$ be a probability distribution. For a partition $M=\{M_{1},\ldots,M_{m}\}$ of the set of nodes of a network, we define the probability of a transition out of community $M_{i}$ with an initial state from the distribution $\pi$ by $q_{i\curvearrowright}\mathrel{\mathop{:}}=\sum_{j\in M_{i},k\notin M_{i}}\pi_{j}p_{kj}$ and the probability of a transition into community $M_{i}$ with an initial state from the distribution $\pi$ by $q_{i\curvearrowleft}\mathrel{\mathop{:}}=\sum_{k\in M_{i},j\notin M_{i}}\pi_{j}p_{kj}$. The probability of a transition into a community that is different from the community of an initial state from the distribution $\pi$ is $q_{\curvearrowleft}\mathrel{\mathop{:}}=\sum_{i\in\{1,\ldots,m\}}q_{i\curvearrowleft}$. The optimal mean encoding length that is associated with the probabilities of transitions into a different community is the entropy [^1] The entropy of a distribution with strictly positive probabilities $p_{1},\ldots,p_{r}$ is ($\mathcal{H}(\{p_{1},\ldots,p_{r}\})\mathrel{\mathop{:}}=-\sum_{i}p_{i}\log_{2}(p_{i})$. $\mathcal{H}(\mathcal{Q})$, where the distribution $\mathcal{Q}$ is
$\mathcal{Q}\mathrel{\mathop{:}}=\left(q_{1\curvearrowleft}/q_{\curvearrowleft},\ldots,q_{m\curvearrowleft}/q_{\curvearrowleft}\right)^{\operatorname{\,T}}\,.$
@@
$\mathcal{P}^{i}\mathrel{\mathop{:}}={\left(q_{i\curvearrowright}/p_{\circlearrowright}^{i},\pi_{k_{1}}/p_{\circlearrowright}^{i},\ldots,\pi_{k_{{n_{i}}}}/p_{\circlearrowright}^{i}\right)^{\operatorname{\,T}}\,,}$
where $M_{i}=\{k_{1},\ldots,k_{{n_{i}}}\}$ (with $k_{1}<\cdots<k_{{n_{i}}}$) and $p_{\circlearrowright}^{i}\mathrel{\mathop{:}}=q_{i\curvearrowright}+\sum_{j\in M_{i}}\pi_{j}$ is a normalization factor. [^2] If $\pi$ is the stationary distribution corresponding to $P$, then $\pi_{k_{1}},\ldots,\pi_{k_{{n_{i}}}}$ are the associated probabilities that a transition ends at a node in community $M_{i}$ and we can interpret the probabilities in $\mathcal{P}^{i}$ as the probabilities of transitions that start in community $M_{i}$ and either leave it or remain in community $M_{i}$. For the partition $M$, the map function that is associated with $P$ and $\pi$ is
$L(M,P,\pi)\mathrel{\mathop{:}}=q_{\curvearrowleft}\mathcal{H}(\mathcal{Q})+\sum_{i=1}^{m}p_{\circlearrowright}^{i}\mathcal{H}(\mathcal{P}^{i})\,.$ (1)
@@
(Map function associated with an adjacency matrix $A$) Let $A$ be an adjacency matrix such that $AW^{-1}$ is regular, where $W\mathrel{\mathop{:}}=\operatorname{diag}\{\omega_{1},\ldots,\omega_{n}\}$ and $\omega_{j}\mathrel{\mathop{:}}=\sum_{i}a_{ij}\neq 0$ for $j\in\{1,\ldots,n\}$ is the out-degree of node $j$. The map function $L(M)$ that is associated with $A$ is
$L(M)\mathrel{\mathop{:}}=L(M,AW^{-1})\,.$
@@ 2.2 Markov time sweeping
The map function (1) is associated with one-step transitions of a random walk. Schaub et al. [39] incorporated Markov time sweeping [^4] Markov sweeping time had been used previously in contexts other than InfoMap [8, 23]. into InfoMap to tune the time scales of transitions by encoding transitions with steps of any length $t>0$. We think of the Markov chain that is determined by $T_{J}\mathrel{\mathop{:}}=AW^{-1}$ (with $AW^{-1}$ specified as in Definition 3) as having time steps of size $1$. Markov time sweeping uses the transition-probability matrix of a continuous-time Markov chain in which the time step is $t$ instead of $1$. We refer to the time $t$ as a “Markov time”. Markov times $t<1$ yield an encoding by the map function at a smaller transition time than $t=1$. This, in turn, yields small communities. By contrast, Markov times $t>1$ yield an encoding at a larger transition time than $t=1$, so the encoding is able to capture transitions of a random walker that take more than one step. We thereby obtain large communities. See [20] for further discussion of encodings and Markov times.
@@ 3 Markov time sweeping and adaptations of InfoMap to absorbing random walks
In this section, we present the adaptations of InfoMap that we use to explore how absorption in a dynamical process can affect community structure. To illustrate the implicit impact of absorption on community structure, consider an absorbing random walk on an undirected and unweighted line network with four nodes (see Figure 1). We represent an absorbing state of an absorbing random walk on a graph as a node [^5] Throughout our paper, our figures do not show the nodes for the absorbing states that are associated with absorbing graphs. Our figures only show nodes that are associated with transient states of the corresponding absorbing random walks. with out-degree 00. For a given node, we refer to the sum of the weights of the edges between the node and absorbing states as the “absorption rate” of the node. Suppose that the absorption rate of node 2 in Figure 1 is much larger than the absorption rates of the other nodes. The large absorption rate of node 2 is a barrier to the absorbing random walk, as transitions from nodes in the set $\{3,4\}$ to nodes in the set $\{1\}$ (and vice versa) are unlikely. The local dynamics (specifically, the absorption at the nodes) induces a partition of the node set that we can interpret as an effective community structure of the network. Such effective community structure can be rather different from the community structure that one detects based on network structure alone.
Figure 1: Consider an absorbing random walk on the depicted four-node network, and suppose that the absorption rate of node 2 is much larger than the absorption rates of the other nodes. The sizes of the nodes and the lengths of the vertical dark-pink segments are positively correlated with the corresponding absorption rates. Detecting communities via modularity maximization or the standard InfoMap algorithm produces a partition of the network into a single community that includes all nodes. However, the flow of an absorbing random walk is trapped in either the set $\{1\}$ (in dark blue) or the set $\{3,4\}$ (in light blue). Consequently, a partition that separates node 1 from nodes 3 and 4 better captures the dynamics of an absorbing random walk than a partition of the network into a single community.
@@
(Absorption-scaled graph) Let $G$ be a directed and weighted graph with adjacency matrix $A=(a_{ij})_{i,j\in\{1,\ldots,n\}}$, where $a_{ij}$ encodes the weight of the edge from node $j$ to node $i$. Let $\vec{d}=(d_{1},\ldots,d_{n})^{\operatorname{\,T}}$ be a vector (which we call an “absorption-rate vector”) with positive entries that we call the “absorption rates”. We define the absorption-scaled graph that is associated with the pair $(G,\vec{d})$ as the graph $\tilde{G}$ with adjacency matrix $\tilde{A}\mathrel{\mathop{:}}=AD^{-1}$, where $D\mathrel{\mathop{:}}=\operatorname{diag}\{d_{1},\ldots,d_{n}\}$.
Figure 2: An absorption-scaled graph. (a) A graph $G$ with absorption-rate vector $\vec{d}$. The pink node has a large absorption rate. (b) The associated absorption-scaled graph $\tilde{G}$, where the arrow length is proportional to the corresponding edge weight. In (a) and (b), node size is proportional to the corresponding absorption rate.
@@
Let $A=(a_{ij})_{i,j\in\{1,\ldots,n\}}$ be the adjacency matrix of a graph $G$. Let $\vec{\delta}=(\delta_{1},\dots,\delta_{n})^{\operatorname{\,T}}$ be a vector with positive entries, where $\delta_{i}$ is the node-absorption rate of node $i$ and is specified independently from the matrix $A$. We refer to the vector $\vec{\delta}$ as the node-absorption-rate vector. Let $D_{\delta}\mathrel{\mathop{:}}=\operatorname{diag}\{\vec{\delta}\}$ and consider a diagonal matrix $H=\operatorname{diag}\{h_{1},\ldots,h_{n}\}$, where $h_{i}\geq 0$. The scaled rate vector $\vec{d}_{\mathrm{s}}(D_{\delta},H)$ is given by the diagonal of $HW+D_{\delta}$.
@@
~$\tilde{\mathcal{L}}(D_{\delta},H)\mathrel{\mathop{:}}=(W-A)(HW+D_{\delta})^{-1}\,.$ (3)
@@
$P_{e}(D_{\delta},H,t)\mathrel{\mathop{:}}=e^{-t\tilde{\mathcal{L}}(D_{\delta},H)}\,.$ (4)
@@
$P_{l}(D_{\delta},H,t)\mathrel{\mathop{:}}=I-t\tilde{\mathcal{L}}(D_{\delta},H)=\left(I-tW(HW+D_{\delta})^{-1}\right)+tA(HW+D_{\delta})^{-1}\,.$ (5)
@@ 4.1.1.1   The distribution $\pi_{\delta}^{(a)}$
where $\vec{t}=N^{\operatorname{\,T}}\vec{1}=(t_{1},\ldots,t_{n})^{\operatorname{\,T}}$, with $\vec{1}\mathrel{\mathop{:}}=(1,\ldots,1)^{\operatorname{\,T}}$, is the vector whose entries give the expected numbers of steps before absorption starting from each non-absorbing state. For each node pair $(i,j)$, the entry $n_{ij}/t_{j}$ of $\hat{N}$ gives the probability that node $i$ is the last node before absorption if we start at node $j$. This probability depends on the node-absorption rates. Given an initial distribution $\pi_{0}$, we obtain the distribution
@@
Suppose that the Markov chain with transition-probability matrix $AW^{-1}$ is regular. Let $\vec{\delta}$ be a node-absorption-rate vector in which all entries are strictly positive, and let $D_{\delta}\mathrel{\mathop{:}}=\operatorname{diag}\{\vec{\delta}\}$. Let $N=(I-A(W+D_{\delta})^{-1})^{-1}$ be the fundamental matrix of the absorbing Markov chain that is associated with $A$ and $\delta_{1},\ldots,\delta_{n}$. Additionally, let $D_{\vec{t}}$ be the diagonal matrix with the column sums of $N$ in its diagonal, and let $\pi$ be the stationary-distribution vector that is associated with $AW^{-1}$.
@@
where $\lVert\Lambda\rVert_{1}<\epsilon$. Let $Q\mathrel{\mathop{:}}=A(W+D_{\delta})^{-1}.$ Because $\lVert Q^{j}\rVert_{1}\leq\lVert Q\rVert_{1}^{j}\leq\lVert AW^{-1}\rVert_{1}^{j}=1$ and $t_{i}\rightarrow\infty$ as $\lVert\vec{\delta}\rVert_{1}\rightarrow 0$, it follows that
@@
whenever $0<\lVert\vec{\delta}\rVert_{1}<\eta_{0}$, where $\lVert\Delta\rVert_{1}<\epsilon$. Let $\Delta^{\prime}\mathrel{\mathop{:}}=\Lambda+\Delta$. Using (13) and (15), we obtain
@@ 4.1.1.2   The transition-probability matrix $P_{\delta}$
$P_{\delta}\mathrel{\mathop{:}}=P_{l}(D_{\delta},I,1)=(I-W(W+D_{\delta})^{-1})+A(W+D_{\delta})^{-1}=D_{\vec{r}}+Q\,,$ (17)
@@
$T_{j}\mathrel{\mathop{:}}=\min\{n\mathrel{\mathop{:}}X_{n}=X_{n-1}\text{ and }X_{0}=j\}\,.$ (19)
@@
(Map function for an absorbing random walk) Let $A$ be the adjacency matrix of an absorbing random walk, and let $\vec{\delta}$ be the walk’s node-absorption-rate vector. Let $M$ be a partition of the node set that is associated with $A$, and let $\pi_{0}$ be an initial probability distribution. We define the map function $L^{(a)}(M,A,\vec{\delta},\pi_{0})\mathrel{\mathop{:}}=L(M,P_{\delta},\hat{N}\pi_{0})$ with $P_{\delta}$ in (17) and $\hat{N}$ by (10).

### v4

@@ 2 Background on InfoMap
$\tilde{A}$ $\begin{pmatrix}A&\vec{0}\\ \vec{\delta}^{\operatorname{\,T}}&0\end{pmatrix}$ Adjacency matrix of a graph with an absorbing state
@@
$Q$ $A(W+D_{\delta})^{-1}$ Submatrix of the transition-probability matrix $\tilde{P}$
for transitions between non-absorbing states
@@ 2.1 The map function and the standard InfoMap algorithm
(Map function) Let $P=(p_{ij})_{i,j\in\{1,\ldots,n\}}$ be a stochastic matrix, and let $\pi=(\pi_{1},\ldots,\pi_{n})^{\operatorname{\,T}}$ be a probability distribution. For a partition $M=\{M_{1},\ldots,M_{m}\}$ of the set of nodes of a network, we define the probability of a transition out of community $M_{i}$ with an initial state from the distribution $\pi$ by $q_{i\curvearrowright}:=\sum_{j\in M_{i},k\notin M_{i}}\pi_{j}p_{kj}$ and the probability of a transition into community $M_{i}$ with an initial state from the distribution $\pi$ by $q_{i\curvearrowleft}:=\sum_{k\in M_{i},j\notin M_{i}}\pi_{j}p_{kj}$. The probability of a transition into a community that is different from the community of an initial state from the distribution $\pi$ is $q_{\curvearrowleft}:=\sum_{i\in\{1,\ldots,m\}}q_{i\curvearrowleft}$. The optimal mean encoding length that is associated with the probabilities of transitions into a different community is the entropy [^1] The entropy of a distribution with strictly positive probabilities $p_{1},\ldots,p_{r}$ is ($\mathcal{H}(\{p_{1},\ldots,p_{r}\}):=-\sum_{i}p_{i}\log_{2}(p_{i})$. $\mathcal{H}(\mathcal{Q})$, where the distribution $\mathcal{Q}$ is
$\mathcal{Q}:=\left(q_{1\curvearrowleft}/q_{\curvearrowleft},\ldots,q_{m\curvearrowleft}/q_{\curvearrowleft}\right)^{\operatorname{\,T}}\,.$
@@
$\mathcal{P}^{i}:={\left(q_{i\curvearrowright}/p_{\circlearrowright}^{i},\pi_{k_{1}}/p_{\circlearrowright}^{i},\ldots,\pi_{k_{{n_{i}}}}/p_{\circlearrowright}^{i}\right)^{\operatorname{\,T}}\,,}$
where $M_{i}=\{k_{1},\ldots,k_{{n_{i}}}\}$ (with $k_{1}<\cdots<k_{{n_{i}}}$) and $p_{\circlearrowright}^{i}:=q_{i\curvearrowright}+\sum_{j\in M_{i}}\pi_{j}$ is a normalization factor. [^2] If $\pi$ is the stationary distribution corresponding to $P$, then $\pi_{k_{1}},\ldots,\pi_{k_{{n_{i}}}}$ are the associated probabilities that a transition ends at a node in community $M_{i}$ and we can interpret the probabilities in $\mathcal{P}^{i}$ as the probabilities of transitions that start in community $M_{i}$ and either leave it or remain in community $M_{i}$. For the partition $M$, the map function that is associated with $P$ and $\pi$ is
$L(M,P,\pi):=q_{\curvearrowleft}\mathcal{H}(\mathcal{Q})+\sum_{i=1}^{m}p_{\circlearrowright}^{i}\mathcal{H}(\mathcal{P}^{i})\,.$ (1)
@@
(Map function associated with an adjacency matrix $A$) Let $A$ be an adjacency matrix such that $AW^{-1}$ is regular, where $W:=\operatorname{diag}\{\omega_{1},\ldots,\omega_{n}\}$ and $\omega_{j}:=\sum_{i}a_{ij}\neq 0$ for $j\in\{1,\ldots,n\}$ is the out-degree of node $j$. The map function $L(M)$ that is associated with $A$ is
$L(M):=L(M,AW^{-1})\,.$
@@ 2.2 Markov time sweeping
The map function (1) is associated with one-step transitions of a random walk. Schaub et al. [39] incorporated Markov time sweeping [^4] Markov sweeping time had been used previously in contexts other than InfoMap [8, 23]. into InfoMap to tune the time scales of transitions by encoding transitions with steps of any length $t>0$. We think of the Markov chain that is determined by $T_{J}:=AW^{-1}$ (with $AW^{-1}$ specified as in Definition 3) as having time steps of size $1$. Markov time sweeping uses the transition-probability matrix of a continuous-time Markov chain in which the time step is $t$ instead of $1$. We refer to the time $t$ as a “Markov time”. Markov times $t<1$ yield an encoding by the map function at a smaller transition time than $t=1$. This, in turn, yields small communities. By contrast, Markov times $t>1$ yield an encoding at a larger transition time than $t=1$, so the encoding is able to capture transitions of a random walker that take more than one step. We thereby obtain large communities. See [20] for further discussion of encodings and Markov times.
@@ 3 Markov time sweeping and adaptations of InfoMap to absorbing random walks
In this section, we present the adaptations of InfoMap that we use to explore how absorption in a dynamical process can affect community structure. To illustrate the implicit impact of absorption on community structure, consider an absorbing random walk on an undirected and unweighted line network with four nodes (see Figure 1). We represent an absorbing state of an absorbing random walk on a graph as a node [^5] Throughout our paper, our figures do not show the nodes for the absorbing states that are associated with absorbing graphs. Our figures only show nodes that are associated with transient (i.e., non-absorbing) states of the corresponding absorbing random walks. with out-degree 00. For a given node, we refer to the sum of the weights of the edges between the node and absorbing states as the “absorption rate” of the node. Suppose that the absorption rate of node 2 in Figure 1 is much larger than the absorption rates of the other nodes. The large absorption rate of node 2 is a barrier to the absorbing random walk, as transitions from nodes in the set $\{3,4\}$ to nodes in the set $\{1\}$ (and vice versa) are unlikely. The local dynamics (specifically, the absorption at the nodes) induces a partition of the node set that we can interpret as an effective community structure of the network. Such effective community structure can be rather different from the community structure that one detects based on network structure alone.
Figure 1: Consider an absorbing random walk on the depicted four-node network, and suppose that the absorption rate of node 2 is much larger than the absorption rates of the other nodes. Detecting communities via modularity maximization or the standard InfoMap algorithm produces a partition of the network into a single community that includes all nodes. However, the flow of an absorbing random walk is trapped in either the set $\{1\}$ (in dark blue) or the set $\{3,4\}$ (in light blue). Consequently, a partition that separates node 1 from nodes 3 and 4 better captures the dynamics of an absorbing random walk than a partition of the network into a single community.
@@
(Absorption-scaled graph) Let $G$ be a directed and weighted graph with adjacency matrix $A=(a_{ij})_{i,j\in\{1,\ldots,n\}}$, where $a_{ij}$ encodes the weight of the edge from node $j$ to node $i$. Let $\vec{d}=(d_{1},\ldots,d_{n})^{\operatorname{\,T}}$ be a vector (which we call an “absorption-rate vector”) with positive entries that we call the “absorption rates”. We define the absorption-scaled graph that is associated with the pair $(G,\vec{d})$ as the graph $\tilde{G}$ with adjacency matrix $\tilde{A}:=AD^{-1}$, where $D:=\operatorname{diag}\{d_{1},\ldots,d_{n}\}$.
Figure 2: An absorption-scaled graph. (a) A graph $G$ with absorption-rate vector $\vec{d}$. The pink node has a large absorption rate. (b) The associated absorption-scaled graph $\tilde{G}$, where the arrow length is proportional to the corresponding edge weight.
@@
Let $A=(a_{ij})_{i,j\in\{1,\ldots,n\}}$ be the adjacency matrix of a graph $G$. Let $\vec{\delta}=(\delta_{1},\dots,\delta_{n})^{\operatorname{\,T}}$ be a vector with positive entries, where $\delta_{i}$ is the node-absorption rate of node $i$ and is specified independently from the matrix $A$. We refer to the vector $\vec{\delta}$ as the node-absorption-rate vector. Let $D_{\delta}:=\operatorname{diag}\{\vec{\delta}\}$ and consider a diagonal matrix $H=\operatorname{diag}\{h_{1},\ldots,h_{n}\}$, where $h_{i}\geq 0$. The scaled rate vector $\vec{d}_{\mathrm{s}}(D_{\delta},H)$ is given by the diagonal of $HW+D_{\delta}$.
@@
~$\tilde{\mathcal{L}}(D_{\delta},H):=(W-A)(HW+D_{\delta})^{-1}\,.$ (3)
@@
$P_{e}(D_{\delta},H,t):=e^{-t\tilde{\mathcal{L}}(D_{\delta},H)}\,.$ (4)
@@
$P_{l}(D_{\delta},H,t):=I-t\tilde{\mathcal{L}}(D_{\delta},H)=\left(I-tW(HW+D_{\delta})^{-1}\right)+tA(HW+D_{\delta})^{-1}\,.$ (5)
@@ 4.1.1.1   The distribution $\pi_{\delta}^{(a)}$
where $\vec{t}=N^{\operatorname{\,T}}\vec{1}=(t_{1},\ldots,t_{n})^{\operatorname{\,T}}$, with $\vec{1}:=(1,\ldots,1)^{\operatorname{\,T}}$, is the vector whose entries give the expected numbers of steps before absorption starting from each non-absorbing state. For each node pair $(i,j)$, the entry $n_{ij}/t_{j}$ of $\hat{N}$ gives the probability that node $i$ is the last node before absorption if we start at node $j$. This probability depends on the node-absorption rates. Given an initial distribution $\pi_{0}$, we obtain the distribution
@@
Suppose that the Markov chain with transition-probability matrix $AW^{-1}$ is regular. Let $\vec{\delta}$ be a node-absorption-rate vector in which all entries are strictly positive, and let $D_{\delta}:=\operatorname{diag}\{\vec{\delta}\}$. Let $N=(I-A(W+D_{\delta})^{-1})^{-1}$ be the fundamental matrix of the absorbing Markov chain that is associated with $A$ and $\delta_{1},\ldots,\delta_{n}$. Additionally, let $D_{\vec{t}}$ be the diagonal matrix with the column sums of $N$ in its diagonal, and let $\pi$ be the stationary-distribution vector that is associated with $AW^{-1}$.
@@
where $\lVert\Lambda\rVert_{1}<\epsilon$. Let $Q:=A(W+D_{\delta})^{-1}.$ Because $\lVert Q^{j}\rVert_{1}\leq\lVert Q\rVert_{1}^{j}\leq\lVert AW^{-1}\rVert_{1}^{j}=1$ and $t_{i}\rightarrow\infty$ as $\lVert\vec{\delta}\rVert_{1}\rightarrow 0$, it follows that
@@
whenever $0<\lVert\vec{\delta}\rVert_{1}<\eta_{0}$, where $\lVert\Delta\rVert_{1}<\epsilon$. Let $\Delta^{\prime}:=\Lambda+\Delta$. Using (13) and (15), we obtain
@@ 4.1.1.2   The transition-probability matrix $P_{\delta}$
$P_{\delta}:=P_{l}(D_{\delta},I,1)=(I-W(W+D_{\delta})^{-1})+A(W+D_{\delta})^{-1}=D_{\vec{r}}+Q\,,$ (17)
@@
$T_{j}:=\min\{n:X_{n}=X_{n-1}\text{ and }X_{0}=j\}\,.$ (19)
@@
(Map function for an absorbing random walk) Let $A$ be the adjacency matrix of an absorbing random walk, and let $\vec{\delta}$ be the walk’s node-absorption-rate vector. Let $M$ be a partition of the node set that is associated with $A$, and let $\pi_{0}$ be an initial probability distribution. We define the map function $L^{(a)}(M,A,\vec{\delta},\pi_{0}):=L(M,P_{\delta},\hat{N}\pi_{0})$ with $P_{\delta}$ in (17) and $\hat{N}$ by (10).